From GlueXWiki
Jump to: navigation, search

February 27, 2014 FDC meeting


  1. Installation status
  2. Engineering (Bill)
  3. Electronics problematic channels (Chris, Nick)
  4. Mini-DAQ status FDC E-log (Beni)
  5. Other


Participants: Elton, Fernando, Dave, Chris, Bill, Beni, and Lubomir. Given the urgency of the FDC testing before the insertion, the agenda was somewhat different.

Testing and Installation

- Beni: The tests done with the fADC125 firmware with 10 and 100 samples (hard coded by Cody) showed the same problem: after several thousand good events, there's a corrupted one (with a wrong length) and after that all are corrupted and there's no way to recover. As for the option to return back to the Gerard's code that was certainly working, the only way to do this is by manually re-programing again all the fADCs. Fernando: Cody, Bryan and Hai are working on this problem and they understand the urgency. Fernando hopes that very soon we will have a working code.

- Lubomir: as of today we started testing the cathode channels with a scope (plan B). It will take about a day per package, so certainly by Tuesday we will be ready. With a scope you can't see all the problems (like frequency of the cosmics, or low amplitude noise) but one can identify the symptoms caused by a bad conductive tape: no signal at all or very small amplitude. So, at least we can identify and fix these. If by the next week we have the fADC working we can test all the channels in one night. In any case we decided to plan inserting the FDC at the end of the next week. If we have the fADCs we will immediately test all the channels after the insertion and by March 17 we should know if there are significant problems that will require to pull the FDC back on the platform.

- There were some discussions initiated by Beni and Fernando, about what will be the consequences if the FDC installation is delayed say by a month. The CDC installation will be delayed, but then the other detectors are not ready for installation. After the meeting Elton discussed this with Tim: we are behind the schedule and more delays will make it even more difficult to do everything that is left in shorter time with the same people.


- By the beginning of the next week we will have a list with problematic connectors and Chris will start fixing them. The difficult part will be to fix the 3M connectors (at the electronics side), but at least Chris can open them and try to identify the issue. If we are sure that it is the 3M connector he can fix them later.


- Most of the discussions were about the cathode scan. All the details are in the E-log entries 142 and 143. The main issues are:

    • Why there are measured points in the middle of the cathode where there is no copper.
    • How to explain the extra gap (by ~250 microns) between the strips exactly in the middle of the cathode. The measurements were done in two days with the interruption exactly in the middle. Bill talked to the company to try to re-create this problem by turning the machine off and on, but right now the machine is busy. Another explanation could be a problem in the art work, in the PCAD or gerber file. Roger promised to look into that.
    • It is not clear why the pitches are different for the different foils. Bill: according to Roger the three foils are cut side by side out of the big roll and the elastic properties at different places might be different. Fernando also explained that the photolithography optics can also be a factor.
    • At the place where the foils are spliced the deviations from the nominal are jumping from strip to strip by up to 100microns.

- In any case using three pitch values (one per foil), the gaps between the foils, plus an extra gap in the middle as free parameters, one can model the strip positions with an average resolution of about 100 microns.

- The fast shut-down of the LV is now working thanks to Dave and Hovanes! (forgot to mention this at the meeting).